OpenStudy (caitlinsanders):

A subway train starting from rest leaves a station with a constant acceleration. At the end of 7.09 s, it is moving at 20.1356 m/s. What is the train’s displacement in the first 4.67231 s of motion? Answer in units of m

OpenStudy (danjs):

Constant acceleration starting from rest. \[v _{i}=0\] \[v _{f}=20.1356 ~~~when~~~t=7.09\] \[x _{0}=0\] \[x _{f}=?\] \[\large a =\frac{ \Delta v }{ \Delta t }=\frac{ 20.1356 }{ 7.09 }=2.84 ~~m/s^2\] There is the acceleration, you can use a kinematic equation to geta function for displacement.. \[\large x _{f}-x _{0}=v _{0}*t + \frac{ 1 }{ 2 }*a*t^2\] use original velocity is zero, and acceleratipon is 2.84, initial position is 0. \[\large x _{f}=\frac{ 1 }{ 2 }*2.84*t^2\]

OpenStudy (danjs):

That is it. You can put in a value for time t and get the displacement.